OpenStudy (anonymous):

I need to isolate r3 R= _______1_________ 1/r1+ 1/r2 +1/r3 Please help and respond as fast as you can. This is due tomorrow.

OpenStudy (phi):

start by flipping both sides

OpenStudy (anonymous):

\[\frac{ 1 }{ R }=\frac{ 1 }{ r _{1} }+\frac{ 1 }{ r _{2} }+\frac{ 1 }{ r _{3} }\] \[\frac{ 1 }{ r _{3} }=\frac{ 1 }{R }-\frac{ 1 }{ r _{1} }-\frac{ 1 }{r _{2} }\]

OpenStudy (anonymous):

Phi, By flipping did you mean to multiply both sides by reciprocal?

OpenStudy (anonymous):

Thank you! It seems much simpler to me now

OpenStudy (anonymous):

yw